OpenStudy (anonymous):

Which of the following equations could be the result of using the comparison method to solve the system shown? x + y = 5 2x + y = 7 ? A. -x - 5 = 7 - 2x B. 5 - x = 7 - 2x C. 5 - x = 2x - 7

OpenStudy (anonymous):

x + y = 5 2x + y = 7 y = -x + 5 y = -2x + 7 -2x + 7 = -x + 5

OpenStudy (anonymous):

y = 5 -x y = 7 - 2x so 5-x = 7 - 2x...............Answer B

OpenStudy (anonymous):

Rewrite that in the appropriate order and you have your answer, just remember to keep the proper signs.

OpenStudy (anonymous):

Or just choose Answer B.

OpenStudy (anonymous):

lol sorry jcpd910

OpenStudy (anonymous):

i didnt see u typing :(

OpenStudy (anonymous):

(ノಥ益ಥ)ノ ┻━┻

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

(no homo) lol

OpenStudy (anonymous):

I'm a girl lol